OpenStudy (ashy98):

Assuming x ≠ 0 and y ≠ 0, what is the quotient of

OpenStudy (ashy98):

\[8x ^{9}y ^{3}+2x ^{6}y ^{2}x8x ^{3}y \] \[2x ^{3}y\]

OpenStudy (e.mccormick):

Do you mean \(\dfrac{8x ^{9}y ^{3}+2x ^{6}y ^{2}x8x ^{3}y}{2x^3y}\)

OpenStudy (ashy98):

yes i did know how to do that lol

OpenStudy (e.mccormick):

\dfrac{top}{bottom}

OpenStudy (e.mccormick):

Is there supposed to be more + or - there? Because the \(2x ^{6}y ^{2}x8x ^{3}y\) looks odd.

OpenStudy (ashy98):

\[8x ^{9} y ^{3}+2 x ^{6} y ^{2}+8x ^{3} y\]

OpenStudy (e.mccormick):

OK. Well, I eould look at what you can facor and cancel.

OpenStudy (e.mccormick):

Here is an example: \(\dfrac{12ab^3+6a^2b^2+2a^3b}{4ab} \implies\) \(\dfrac{2(6ab^3+3a^2b^2+a^3b)}{2(2ab)} \implies\) \(\dfrac{\cancel{2}(6ab^3+3a^2b^2+a^3b)}{\cancel{2}(2ab)} \implies\) But that is not all that factors and cancels in this case. I can do an a and a b too. \(\dfrac{ab(6b^2+3ab+a^2)}{2ab} \implies\) \(\dfrac{\cancel{ab}(6b^2+3ab+a^2)}{2\cancel{ab}} \implies\) \(\dfrac{6b^2+3ab+a^2}{2} \)